OpenStudy (mathhater):

Give the 5th and 10th term for the sequence formula below. tn = 1 + 6n

Parth (parthkohli):

Substitute n with 5 to get the 5th term. Substitute n with 10 to get the 10th term.

Parth (parthkohli):

\( \color{Black}{\Rightarrow t_5 = 1 + 6(5) }\) \( \color{Black}{\Rightarrow t_{10} = 1 + 6(10) }\)

OpenStudy (anonymous):

t5 = 1 + 6(5) and, t10 = 1 + 6(10) Solve them and find the terms...